Periodic Inventory by Three Methods; Cost of Merchandise Sold The units of an item available for sale during the year were as follows: Jan. 1 Inventory 30 units @ $112 Mar. 10 Purchase 70 units @ $122 Aug. 30 Purchase 30 units @ $126 Dec. 12 Purchase 70 units @ $132 There are 80 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost and the cost of merchandise sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar. Cost of Merchandise Inventory and Cost of Merchandise Sold Inventory Method Merchandise Inventory Merchandise Sold First-in, first-out (FIFO) $ $ Last-in, first-out (LIFO) Weighted average cost
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 30 112 3360
Purchase 1 70 122 8540
Purchase 2 30 126 3780
Purchase 3 70 132 9240
Total 200 24920

Calculate following :

Cost of ending inventory Cost of goods sold
FIFO (70*132+10*126) = 10500 24920-10500 = 14420
LIFO (30*112+50*122) = 9460 24920-9460 = 15460
Weighted average 24920/200*80 = 9968 24920-9968 = 14952